OpenStudy (anonymous):

find two points on the line to be graphed and enter their coordinates in the boxes below A=( , ) B= ( , ) y =5/2x+1

OpenStudy (texaschic101):

you can find the x and y intercepts from the original equation. y = 5/2x + 1 to find x intercept, sub in 0 for y 0 = 5/2x + 1 -5/2x = 1 x = 1/(-5/2) x = 1 * -2/5 x = -2/5 or - 0.4 so your x intercept is (-0.4, 0) to find your y intercept, sub in 0 for x y = 5/2(0) + 1 y = 1 so your y intercept is (0,1) that is your points... A = (-0.4,0) and B = (0,1)